solve this equation by using elimination
3x-2y=10
4x-3y=15
I think you can multiply the top equation by 3 but from there I am not sure

Answer by stanbon(75887)

solve this equation by using elimination
3x-2y=10
4x-3y=15
--------------
To eliminate the "y terms", multiply the
1st equation by 3 and the 2nd equation by 2 to get:
-----
9x - 6y = 30
8x - 6y = 30
-----------------
Subtract the 2nd from the 1st to get:
x = 0
----
Substitute into 3x-2y=10 and solve for "y":
0-2y=10
y = -5
=============================================
Cheers,
Stan H.